### v1

# An Accelerated Hybrid Low-Rank Augmented Lagrangian Method for Large-Scale Semidefinite Programming on GPU

Abstract
This paper introduces cuHALLaR, a GPU-accelerated implementation of the HALLaR method proposed in [47] for solving large-scale semidefinite programming (SDP) problems. We demonstrate how our Julia-based implementation efficiently uses GPU parallelism through optimization of simple, but key, operations, including linear maps, adjoints, and gradient evaluations. Extensive numerical experiments across three problem classes—maximum stable set, matrix completion, and phase retrieval show significant performance improvements over both CPU implementations and existing GPU-based solvers. For the largest instances, cuHALLaR achieves speedups of 30-140x on matrix completion problems, up to 135x on maximum stable set problems for Hamming graphs with 8.4 million vertices, and 15-47x on phase retrieval problems with dimensions up to 3.2 million. Our approach efficiently handles massive problems with dimensions up to $(n,m)\approx(8\times 10^{6},3\times 10^{8})$ at $10^{-5}$ relative precision, solving matrix completion instances with over 8 million rows and columns in just 142 seconds. These results establish cuHALLaR as a very promising GPU-based method for solving large-scale semidefinite programs.
Keywords: semidefinite programming, augmented Lagrangian, low-rank methods, gpu acceleration, Frank-Wolfe method
1 Introduction
Let $\mathbb{S}^{n}$ be the set of $n\times n$ symmetric matrices. The notation $A\succeq B$ means that $A-B$ is positive semidefinite. We are interested in solving the primal-dual pair of semidefinite programs (SDPs)
$P_{*}:=\quad\min_{X}\quad\{C\bullet X\quad:\quad\mathcal{A}(X)=b,\quad X\in\Delta^{n}\}$ (P)
and
$D_{*}:=\quad\max_{p\in\mathbb{R}^{m},\theta\in\mathbb{R}}\quad\{-b^{\top}p-\theta\quad:\quad S\vcentcolon=C+\mathcal{A}^{*}(p)+\theta I\succeq 0,\quad\theta\geq 0\}$ (D)
where $b\in\mathbb{R}^{m}$, $C\in\mathbb{S}^{n}$, $\mathcal{A}:\mathbb{S}^{n}\to\mathbb{R}^{m}$ is a linear map, $\mathcal{A}^{*}:\mathbb{R}^{m}\to\mathbb{S}^{n}$ is its adjoint, and $\Delta^{n}$ is the spectraplex
$\Delta^{n}\vcentcolon=\{X\in\mathbb{S}^{n}:\mathrm{Tr}(X)\leq 1,X\succeq 0\}.$ (1)
Semidefinite programming arises across a wide spectrum of applications, notably in machine learning, computational sciences, and various engineering disciplines. Despite its versatility, efficiently solving large-scale SDPs remains computationally demanding. Interior-point methods, although widely employed, are often constrained by computational complexity as they often rely on solving a large linear system using a Newton-type method, which involves an expensive inversion of a large matrix [5, 2, 31, 46, 65]. These computational bottlenecks have sparked interest in developing efficient first-order methods for solving large-scale problems. Unlike interior-point methods, first-order methods scale well with the dimension of the problem instance and predominantly utilize sparse matrix-vector multiplications to perform function and gradient evaluations. For an extensive survey and analysis of recent developments in using first-order methods for solving SDPs, we refer the reader to the comprehensive literature available in [1, 16, 17, 19, 20, 23, 25, 33, 26, 38, 34, 51, 54, 56, 58, 71, 74, 76, 78, 81, 82].
Due to GPU’s great capabilities in parallelizing matrix-vector multiplications, they have been shown to work extremely well in speeding up first-order methods in practice [12, 30, 39, 43, 42, 41, 57]. In this paper, we introduce an enhanced Julia-based GPU implementation, referred to as cuHALLaR, of the first-order HALLaR method developed in [47] for solving huge-scale SDPs. Our work reinforces the viability of using GPU architectures for efficient parallelization in semidefinite programming solvers. Moreover, our experiments show potential for excellent speedups when using GPUs for optimization solvers.
1.1 Related literature
We divide our discussion into four parts: first, the emerging literature on utilizing GPUs for mathematical optimization, second, the challenges in solving large SDPs, third, recent techniques in solving large SDPs using low-rank approaches, and fourth, a comparison between cuHALLaR, a recent GPU SDP solver cuLoRads, and the CPU version of HALLaR.
GPUs for Mathematical Programming
GPUs offer massive parallelism that accelerates mathematical programming through parallel linear algebra operations. Recent GPU-based solvers for linear programming and quadratic programming [41, 43, 42] have achieved performance which outperforms traditional CPU-based methods. First-order methods map well to GPU architectures, unlike simplex or interior-point methods which rely on difficult-to-parallelize factorizations. For linear programming, cuPDLP.jl incorporates GPU-accelerated sparse matrix operations to match commercial solver performance. Quadratic programming solvers like cuOSQP [57] and CuClarabel [12] similarly exploit GPU parallelism for significant speedups. Most recently, [39] proposed a new GPU conic programming solver that utilizes the primal-dual hybrid gradient (PDHG) method for solving several classes of cone optimization problems, including semidefinite programs.
Challenges in large-scale semidefinite programming
Solving large-scale SDPs in practice presents greater challenges compared to solving large-scale linear programs and quadratic programs due to the reliance of many SDP solvers on full eigendecomposition, inversion, and storage of large dense matrices. To circumvent several of these issues, pre-processing techniques such as facial reduction, chordal conversion, and symmetry reduction have been proposed which can potentially convert the original SDP (P) into a smaller one or one with a more favorable structure [66, 24, 27, 80, 83, 79, 32, 6, 21, 61, 53]. Interior-point methods are then often applied to the smaller reformulated SDP. Another popular approach that is commonly used in practice to solve large-scale SDPs and which avoids storing and inverting large dense matrices is the low-rank approach first proposed by Burer and Monteiro [9, 10]. The next paragraph describes low-rank first-order methods for solving SDPs in more detail.
Low-rank first-order methods for SDPs
The low-rank approach proposed by Burer and Monteiro [9, 10] converts SDP (P) into a nonconvex quadratic program with quadratic constraints (QCQP) through the transformation $X=UU^{T}$, where $U\in\mathbb{R}^{n\times r}$ and $r\ll n$. The main advantage of this approach is that the QCQP has $n\times r$ variables, while (P) has $n(n+1)/2$ variables. Also, the PSD constraint $X\succeq 0$ has been eliminated through the low-rank transformation so methods do not need to perform eigendecompositions of dense matrices to approximately solve the QCQP. Burer and Monteiro proposed an efficient first-order limited-memory BFGS augmented Lagrangian (AL) method, which only relied on matrix-vector multiplications and storage of $n\times r$ matrices, to find an approximate local minimizer of the QCQP. Recent landscape results have shown that if the factorization parameter $r$ is taken large enough, then finding local minimizers of the QCQP is actually equivalent to finding global minimizers of the SDP (P) [14, 15, 7, 8, 4, 55, 67]. For a more extensive list of papers which characterize the optimization landscape of nonconvex formulations obtained through Burer-Monteiro factorizations, see [50, 22, 44, 40, 49, 18, 45, 37]. For recent advances and developments in accelerating low-rank first-order methods for solving SDPs, we refer the reader to the extensive literature available in [23, 29, 30, 47, 62, 63, 64, 69, 70, 77, 68, 76, 78].
We now give a brief comparison between HALLaR, which was first developed in [47] and the LoRADS method developed in [29]. HALLaR utilizes an inexact AL method to solve (P). It solves each of its AL subproblems over the spectraplex $\operatorname*{arg\,min}_{X\in\Delta^{n}}\quad C\bullet X+p^{\top}(\mathcal{A}(X)-b)+\frac{\beta}{2}\|\mathcal{A}X-b\|^{2}$ by first restricting them to the space of matrices of rank at most $r$ through the transformation $X=UU^{T}$ and then applying a nonconvex solver to the reformulation. It uses a low-rank Frank-Wolfe step to escape from a possible spurious stationary point obtained by the nonconvex solver, a step which also potentially increases the rank of the current iterate.
LoRADS, on the other hand, adopts a two-stage approach. In their first stage, they utilize the low-rank limited-memory BFGS AL method of Burer and Monteiro as a warm-start strategy for their second stage. The second stage of LoRADS also utilizes an inexact AL method applied to (P) but the major difference being the way the AL subproblems are reformulated. In contrast to the $X=UU^{T}$ change of variables used by HALLaR, LoRADS reformulates them by replacing $X$ by $UV^{T}$ and adding the constraint $U=V$. It then approximately solves these nonconvex reformulations of the AL subproblems by an alternating minimization penalty-type approach which penalizes $U=V$ constraint and alternately minimizes with respect to the blocks $U$ and $V$ using the conjugate gradient method. Similar to HALLaR, the inexact stationary solutions of the nonconvex AL reformulations may not be near global solutions of the AL subproblems as they are not necessarily equivalent. In contrast to HALLaR, LoRADS uses a heuristic that increases the number of columns of $U$ and $V$ to attempt to escape from the these spurious solutions.
cuHALLaR versus cuLoRADS and HALLaR
Like cuLoRads, which is a GPU accelerated version of LoRads, the cuHALLaR method proposed in this paper is a GPU accelerated version of HALLaR. It is displayed on several important classes of SDPs that cuHALLaR is 2 to 25 times faster than cuLoRads on many large instances. cuHALLaR can efficiently solve massive problems to $10^{-5}$ relative precision in just a few minutes. For example, cuHALLaR is able to solve a matrix completion SDP instance where the size of the matrix variable is 8 million and the number of constraints is approximately 300 million in just 142 seconds.
cuHALLaR also achieves massive speedup compared to its CPU counterpart, HALLaR. cuHALLaR achieves speedups of 30-140x on large matrix completion SDP instances, up to 135x speedup on large maximum stable set SDP instances, and 15-47x speedup on phase retrieval SDP instances. These numerical results show that cuHALLaR is a very promising GPU-based method for solving extremely large SDP instances.
1.2 Notation
[…]
For a given $\epsilon\geq 0$, the $\epsilon$-normal cone of a closed convex set $C$ at $z\in C$, denoted by $N^{\epsilon}_{C}(z)$, is
$N^{\epsilon}_{C}(z):=\{\xi\in\mathbb{R}^{n}:\langle\xi,u-z\rangle\leq\epsilon,\quad\forall u\in C\}.$
The normal cone of a closed convex set $C$ at $z\in C$ is denoted by $N_{C}(z)=N^{0}_{C}(z)$. Finally, we define the Frobenius ball of radius $r$ in $\mathbb{R}^{n\times s}$ space to be
$B_{r}^{s}:=\{U\in\mathbb{R}^{n\times s}:\|U\|_{F}\leq r\}.$ (2)
2 Overview of HALLaR
This section provides a minimal review of HALLaR, the hybrid low-rank augmented Lagrangian method that was first developed in [47].
As mentioned in the Introduction, HALLaR is an inexact AL method that approximately solves the pair of SDPs (P) and (D). Given a tolerance pair $(\epsilon_{p},\epsilon_{pd})\in\mathbb{R}^{2}_{++}$, HALLaR aims to find a $(\epsilon_{p},\epsilon_{pd})$-optimal solution of (P) and (D), i.e., a triple $(\bar{X},\bar{p},\bar{\theta})\in\Delta^{n}\times\mathbb{R}^{m}\times\mathbb{R}_{+}$ that satisfies
$\|\mathcal{A}\bar{X}-b\|\leq\epsilon_{p},\quad|C\bullet\bar{X}-(b^{T}\bar{p}-\bar{\theta})|\leq\epsilon_{pd},\quad C+\mathcal{A}^{*}p+\bar{\theta}I\succeq 0,\quad\bar{\theta}\geq 0.$ (3)
Being an inexact AL method, HALLaR generates sequences $\{X_{t}\}$ and $\{p_{t}\}$ according to the following updates
$X_{t}\quad$ $\approx\quad\operatorname*{arg\,min}_{X}\quad\{\mathcal{L}_{\beta_{t}}(X;p_{t-1})\;:\;X\in\Delta^{n}\},$ (4a)
$p_{t}\quad$ $=\quad p_{t-1}+\beta_{t}(\mathcal{A}(X_{t})-b)$ (4b)
where
$\mathcal{L}_{\beta}(X;p)\quad\vcentcolon=\quad C\bullet X+p^{\top}(\mathcal{A}(X)-b)+\frac{\beta}{2}\|\mathcal{A}X-b\|^{2}$ (5)
is the AL function. The key part of HALLaR is an efficient method, called the hybrid low-rank method (HLR), for finding a near-optimal solution $X_{t}$ of the AL subproblem (4a). HLR never forms $X_{t}$ but outputs a low-rank factor $U_{t}\in B_{1}^{s_{t}}$ such that $X_{t}=U_{t}U_{t}^{T}$ where $B_{1}^{s_{t}}$ is as in (2) and hopefully $s_{t}\ll n$. The next paragraph briefly describes HLR in more detail for solving (4a).
Given a pair $(s,\tilde{Y})\in\mathbb{Z}_{+}\times B_{1}^{s}$, which is initially set to $(s,\tilde{Y})=(s_{t-1},U_{t-1})$, a general iteration of HLR performs the following steps: i) it computes a suitable stationary point $Y$ of the nonconvex reformulation of (4a) obtained through the change of variable $X=UU^{T}$, namely:
$\min_{U}\quad\left\{g_{t}(U)=\mathcal{L}_{\beta_{t}}(UU^{\top};p_{t-1})\quad:\quad U\in B_{1}^{s}\right\};$ ($\mathrm{L}_{r}$)
(ii) checks whether $X=YY^{\top}$, is a reasonable approximate solution of (4a); if so it stops and sets $(s_{t},U_{t})=(s,Y)$; and; iii) else, it performs a Frank-Wolfe step in the $X$-space (but implemented in the $U$-space) to obtain a new pair $(s,\tilde{Y})$.
We now give more details about the above steps. First, we consider step i). This step is implemented with the aid of a generic nonlinear solver which, started from $\tilde{Y}$, computes a $(\rho;\tilde{Y})$-stationary solution of ($\mathrm{L}_{r}$), i.e., a triple $(Y;R,\rho)$ such that
$R\in\nabla g_{t}(Y)+N_{B^{s}_{1}}(Y),\quad\|R\|_{F}\leq\rho,\quad g_{t}(Y)\leq g_{t}(\tilde{Y}).$ (6)
It is easy to verify that (6), in terms of the SDP data, is equivalent to
$R\in 2[C+\mathcal{A}^{*}(p_{t-1}+\beta_{t}(\mathcal{A}(YY^{T})-b))]Y+N_{B^{s}_{1}}(Y),\quad\|R\|_{F}\leq\rho.$
Appendix A describes a method that is able to carry out step i), namely, ADAP-AIPP, whose exact formulation and analysis can be found in [47, 60], and also in earlier works (see e.g. [11, 52, 35, 36]) in other related forms.
The details of steps ii) and iii) implemented in the $X$-space are quite standard. The details of its implementation in the $U$-space are described in the formal description of the algorithm given below. The only point worth observing at this stage is that, due to the rank-one update nature of the Frank-Wolfe method, at the end of step iii), the low-rank parameter $s$ is set to either $s=1$ (unlikely case) or $s=s+1$ (usual case). Hence, in the usual case, the number of columns of $\tilde{Y}$ increases by one.
We now formally state HALLaR below.
Algorithm 1 HALLaR. Output: $(\bar{X},\bar{p},\bar{\theta})\in\Delta^{n}\times\mathbb{R}^{m}\times\mathbb{R}_{+}$, an $(\epsilon_{\mathrm{p}},\epsilon_{\mathrm{pd}})$-solution of the pair of SDPs (P) and (D).
1:Let initial points $(U_{0},p_{0})\in B^{s_{0}}_{1}\times\mathbb{R}^{m}$, and tolerance pair $(\epsilon_{\mathrm{p}},\epsilon_{\mathrm{pd}})\in\mathbb{R}^{2}_{++}$ be given, and let $\left\{\epsilon_{t}\right\}_{t\geq 1}$ and $\left\{\beta_{t}\right\}_{t\geq 1}$ be a decreasing and increasing sequence of positive integers, respectively.
2:set $t\leftarrow 1$;
3:set $\tilde{Y}=U_{t-1}$, $s=s_{t-1}$, $Y=0_{n\times s}$, $G=0_{n\times n}$, and $\theta=\infty$;
[…]
5:     call a nonconvex solver with initial point $\tilde{Y}$, tolerance $\epsilon_{t}$, and function $\mathcal{L}_{\beta_{t}}(UU^{\top};p_{t-1})+\delta_{B_{1}^{s}}(U)$ to find a point $Y\in B^{s}_{1}$ that is an $\epsilon_{t}$-approximate stationary solution (according to the criterion in (6)) of
$\min_{U}\quad\left\{\mathcal{L}_{\beta_{t}}(UU^{\top};p_{t-1})\quad:\quad\|U\|_{F}\leq 1,\quad U\in\mathbb{R}^{n\times s}\right\};$ (7)
6:     compute
$G:=\nabla(\mathcal{L}_{\beta_{t}}(YY^{T};p_{t-1}))\in\mathbb{S}_{n}$ (8)
and a minimum eigenpair of $G$, i.e., $(\lambda_{\min}(G),v_{\min}(G))\in\mathbb{R}\times\mathbb{R}^{n}$, and set
$\theta:=\max\{-\lambda_{\min}(G),0\},\qquad y:=\begin{cases}v_{\min}(G)&\text{if $\theta>0$}\\ 0&\text{ otherwise};\end{cases}$ (9)
7:     if $(GY)\bullet Y+\theta\leq\epsilon_{t}$ then
8:         break while loop and go to step 11 below;
9:     end if
10:     compute
[…]
and set
$(\tilde{Y},s)=\begin{cases}(y,1)&\text{ if $\alpha=1$}\\ \left(\left[\sqrt{1-\alpha}\,Y,\sqrt{\alpha}\,{y}\right],s+1\right)&\text{ otherwise};\end{cases}$ (11)
11:end while
12:set $(U_{t},\theta_{t},s_{t})=(Y,\theta,s)$;
13:set
$p_{t}=p_{t-1}+\beta(\mathcal{A}(U_{t}U_{t}^{\top})-b);$ (12)
14:if
$\|\mathcal{A}(U_{t}U_{t}^{\top})-b\|\leq\epsilon_{\mathrm{p}},\quad|CU_{t}\bullet U_{t}-(-b^{\top}p_{t}-\theta_{t})|\leq\epsilon_{pd}$ (13)
then return $(\bar{U},\bar{p}_{,}\bar{\theta})=(U_{t},p_{t},\theta_{t})$;
15:set $t=t+1$ and go to step 2.
Several remarks about each of the steps of HALLaR are now given. First, $t$ is the iteration count for HALLaR. Second, each iteration of HALLaR performs a HLR call and each call is performed in the while loop from steps 3 to 10. In line 2, the triple $(Y,G,\theta)$ is set as $(0_{n\times s},0_{n\times n},\infty)$, which means the HLR call in the while loop is always performed during each iteration of HALLaR. Third, after a HLR call is finished, HALLaR updates the Lagrange multiplier in line 12 and checks for termination in line 13. Lastly, the termination condition on this line does not check for near dual feasibility since it can be shown that for every $t\geq 1$,
$S_{t}:=C+\mathcal{A}^{*}p_{t}+\theta_{t}I\succeq 0,\quad\theta_{t}\geq 0,$ (14)
and hence that HALLaR always generates feasible dual iterates $(p_{t},\theta_{t},S_{t})$’s.
We now comment on the steps performed within the while loop (lines 3 to 10) in light of the three steps of the HALLaR outline given just before its formal description. Step 4 is implementing step i) of the HALLaR outline. Steps 5 and 6 are implementing step ii). Finally, step 9 implements step iii).
We now give an interpretation of step 9 of HALLaR relative to the $X$-method. The stepsize $\alpha$ computed in (10) corresponds to solving
$\operatorname*{arg\,min}_{\alpha}\left\{\mathcal{L}_{\beta_{t}}\left(\alpha yy^{T}+(1-\alpha)YY^{T};p_{t-1}\right):\alpha\in[0,1]\right\}\}.$
Moreover, the iterate $\tilde{Y}$ computed in equation (11) has the property that $\tilde{Y}\tilde{Y}^{\top}$ is equal to the usual FW iterate $\alpha yy^{T}+(1-\alpha)YY^{T}$ in the $X$-space; hence $\tilde{Y}$ is an $U$-factor for the Frank-Wolfe iterate in the $X$-space.
3 Accelerating Hallar
In this section, we discuss how to translate the initial CPU implementation of HALLaR efficiently to a scalable and well-optimized GPU implementation.
3.1 GPU architecture, threads, and minimizing CPU/GPU communication
GPUs execute threads in a Single Instruction, Multiple Data (SIMD) fashion: groups of 32 threads, called warps, run the same instruction in lockstep. These warps are organized into blocks, each mapped onto a streaming multiprocessor (SM). Within a block, threads can coordinate via barrier synchronizations (e.g. “__syncthreads”) and share data through fast on‑chip shared memory. Kernels (functions designed to run in parallel in the GPU) launch a grid of blocks to cover the entire data set, so that thousands of warps can run concurrently, hiding memory latency and maximizing arithmetic throughput.
The GPU memory hierarchy further shapes performance. Global memory, while large, has high latency and must be accessed coalesced to achieve high bandwidth. Shared memory resides on each SM and offers very low latency when threads within a block reuse it. Finally, each thread has its own registers for the fastest access; spilling to local memory incurs significant cost. Effective GPU kernels minimize global loads/stores, exploit shared memory for data reuse, and balance register usage to avoid spilling.
Figure 1, adapted from the NVIDIA documentation [48], displays the NVIDIA Hopper architecture used in modern GPUs like the H200. The SMs are the main units of execution inside the GPU. They are grouped into the TPC (Texture Processing Clusters) which are in turn grouped into GPC (GPU Processing Clusters). This hierarchy of clusters allows the GPU to improve the management and distribution of the workload between the SMs, improving scalability. The number of available GPCs and TPCs depends on the GPU model. Inside each SM, the warp scheduler, dispatch unit and the L0 instruction cache work as a control unit, delivering instructions to the warps. Next, we see the set of registers each SM has access to, followed by execution components: CUDA cores, tensor cores, LD/ST and SFU. The CUDA cores execute scalar aritmetic operations, while the tensor cores are specialized to execute multiplication and addition $C=D*A+C$ in batches of data. The LD/ST units performs load and store operations on the data. The SFU (Special Function Units) are specialized for transcendental and trigonometric operations. The TMA (Tensor Memory Acceleration) provides asynchronous transfers of large blocks of data between the global and shared memory. Next, we see the L1-cache that optimizes transfers between the shared memory and the CUDA/tensor cores. Back to the GPU, we have the L2-cache that optimizes transfers from the global GPU memory to all SMs. Access to L2-cache by the SMs is controlled by the GPC.
Figure 1: GPU architecture, adapted from the NVIDIA documentation [48], displaying the general GPU design and the components of each SM.
Before launching a kernel, we must transfer the input dataset from the CPU (also known as host in GPU programming) to the GPU (also called device) and define the number of blocks and threads used. The memory allocated in the CPU is not accessible by the GPU and vice-versa. When the kernel is launched, the workload is distributed among the GPCs that, in turn, distribute the workload among their TPCs. The TPCs forward the thread blocks among its SMs. The SMs divide the thread blocks into warps and prepare the shared memory and registers. The warp scheduler and the dispatch unit select the warps ready to run and dispatch the instructions to execution by the CUDA/tensor cores, the LD/ST or the SFU. After the kernel is finished, a transfer from the device back to the host may be necessary.
The GPU programming paradigm is meant to integrate CPU and GPU to obtain the performance gains of both architectures. Sections of the code that allow high parallelization, like matrix, vector and tensor operations are implemented as kernels, while serial tasks are implemented as regular functions in the CPU. This CPU/GPU integration sometimes requires data transfers between the CPU and GPU. Modern GPUs compute faster than they can feed data from memory, so these transfers are often the bottleneck of applications. Therefore we must minimize data transfers to achieve maximum performance.
cuHALLaR stores on the GPU, at the beginning, key data such as the primal/dual solutions, the right-hand side vector $b$, and the data used to compute $C$ and $\mathcal{A}$. Any other auxiliary data is allocated directly on the GPU. So all operations that involve arrays with huge size are done in the GPU, including the dual feasibility computations, which was computed in the CPU by other GPU solvers like [30].
Another important performance aspect is that, just like in the CPU, launching threads in the GPU incurs an overhead due to the grid, context and communication management. Therefore it is also important to minimize the number of kernels launched, merging trivial kernels and avoiding to launch them inside loops when possible.
In HALLaR, three operations are critical to performance: the linear map $\mathcal{A}$, the adjoint $\mathcal{A}^{*}$ and the gradient $\nabla g$. These operations are performed at every ADAP-FISTA iteration, and the gradient is also computed at the Frank-Wolfe step of HLR. For the maximum stable set and matrix completion problems, CuHallar solves $\mathcal{A}$ and $\mathcal{A}^{*}$ in a single kernel, and the gradient is solved with two kernels. In the phase retrieval problem, due to the computation of the discrete Fourier transform, all operations are solved with $r$ kernel calls, where $r$ is the rank of the current solution.
In cuHALLaR, we use the CUDA.jl library, the Julia interface for the CUDA runtime library. CUDA.jl provides high level implementations of many cuBLAS, cuSPARSE and cuFFT functions. The transfer of data between the CPU and GPU is as simple as converting a variable between the Julia Array type and the CUDA.jl CuArray type. CUDA.jl also provides macros so users can implement their own kernels. With the help of Julia broadcast fusion feature, CUDA.jl is able to merge kernels in broadcasted expressions. This allows cuHallar to reduce the number of kernels invoked, specially in the ADAP-FISTA method.
4 Numerical Experiments
In this section, we provide provide extensive experiments comparing cuHALLaR, HALLaR, and cuLoRADS. The experiments concern SDP instances for the matrix completion, the maximum stable set, and the phase retrieval problems. These problems were also used in the benchmark of the original HALLaR paper [47].
4.1 Experimental Setup
Hardware and software.
Our experiments compare the GPU methods cuHALLaR and cuLoRADS, and the CPU method HALLaR. The GPU experiments are performed on an NVIDIA H200 with 142 GB VRAM deployed on a cluster with an Intel Xeon Platinum 8469C CPU. For CPU benchmarks, we use a single Dual Intel Xeon Gold 6226 CPUs @ 2.7 GHz (24 cores/node). Further details are given in Table 1. We set the memory of the CPU benchmarks at 142GB, which is equal to the maximum memory of the GPU.
We implemented cuHALLaR as a Julia [3] module. Our implementation uses the CUDA language [13] for interfacing with NVIDIA CUDA GPUs. The experiments are performed running CUDA 12.6.1 and Julia 1.11.3. The reported running times do not take into account the pre-compilation time.
Specification CPU Node GPU
Processor(s) Dual Intel Xeon Gold 6226 @ 2.70 GHz NVIDIA H200
Cores / SMs 24 cores 80 SMs
Cache (per CPU) L3: 19.25 MB
 (per core) L2: 1 MB; L1: 32 KB (D+I) L1/Shared (per SM), L2
Peak FP64 Perf. $\sim$2.07 TFLOPS $\sim$34 TFLOPS
Memory Size 142 GB DDR4 142 GB HBM3e
Memory Bandwidth $\sim$140.7 GB/s 4.8 TB/s
Table 1: Comparison of CPU and GPU specifications used in experiments.
Termination criteria.
Given a tolerance $\epsilon>0$, all methods stop when an iterate $(X,p)\in\Delta^{n}\times\mathbb{R}^{m}$ satisfying
$\max\left\{\frac{\|\mathcal{A}X-b\|_{2}}{1+\|b\|_{1}},\frac{|\mathrm{pval}-\mathrm{dval}|}{1+|\mathrm{pval}|+|\mathrm{dval}|},\frac{\max\{0,-\lambda_{\min}(C+{\mathcal{A}}^{*}{p})\}}{1+\|C\|_{1}}\right\}\leq\epsilon.$ (15)
is obtained, where $\mathrm{pval}=C\bullet X$ and $\mathrm{dval}=-b^{\top}p$ denote the primal and dual values of $(X,p)$, respectively.
Scaling.
We have assumed in Section 1 that the trace of $X$ is bounded by one. However, cuHALLaR and HALLaR allow $\operatorname*{tr}X\leq 1$ to be replaced by the more general constraint $\operatorname*{tr}(X)\leq\tau$ for some $\tau>0$. This is done by reformulating the SDP as in (P) using the changes of variables $X\leftarrow X/\tau$ and changing the rhs as $b\leftarrow b/\tau$. (15) is used in the experiments of phase retrieval and matrix completion.
Initialization.
HALLaR and cuHALLaR are initialized using an initial random $n\times 1$ matrix with entries generated independently from the Gaussian distribution over $\mathbb{R}$ (or over $\mathbb{C}$ for complex-valued SDPs) which is then scaled to be a unit vector. The dual initial point $p_{0}$ is taken to be the zero vector.
Input format.
HALLaR and cuHALLaR require the following user inputs to describe the SDP instance:
  1. (i)
    The vector $b\in\mathbb{R}^{m}$ and the scalar $\tau>0$. Here, $\tau$ represents the upper bound for the trace of the primal matrix variable $X\in\mathbb{S}^{n}_{+}$, such that $\mathrm{Tr}(X)\leq\tau$.
  2. (ii)
    A routine for evaluating $CU$ for arbitrary matrices $U$ of compatible dimensions.
  3. (iii)
    A routine for evaluating $(\mathcal{A}^{*}p)U$ for arbitrary vectors $p$ and matrices $U$ of compatible dimensions.
  4. (iv)
    A routine for evaluating $\mathcal{A}(UU^{\top})$ for arbitrary matrices $U$ of compatible dimensions.
We are aware that the above operator-based input format for SDPs is not prevalent as matrix-based representations such as the SDPA format [72, 73] or the SEDUMI format [59]. Matrix-based representations have been widely adopted by numerous SDP solvers, and are quite effective for SDPs involving sparse matrices. However, these representations might not be effective for SDPs involving dense matrices. In particular, the stable set and the phase retrieval SDPs involve fully dense matrices, and any code that requires input in a matrix-based format (e.g., SDPA format) can take a long time to read and manipulate their data. In contrast, our operator-based input format enables us to efficiently handle extremely large stable set and phase retrieval SDP instances.
Ratio:
In all tables below, the entry labeled “Ratio” reports the running time of the CPU implementation divided by the running time of the GPU implementation (or, when comparing two solvers, the time of the first solver over the time of the second). For instance, in Table 2 the first row shows cuHALLaR runtime of 0.77 sec and a cuLoRADS runtime of 123.71 sec, yielding a ratio of 160.662. This means that cuLoRADS took 160.662 times longer to solve that instance than cuHALLaR.
4.2 Experiments for matrix completion
Given integers $n_{2}\geq n_{1}\geq 1$, consider the problem of retrieving a low rank matrix $M\in\mathbb{R}^{n_{1}\times n_{2}}$ by observing a subset $\{M_{ij}:(i,j)\in\Omega\}$ of its entries. A standard approach to tackle this problem is by considering the nuclear norm relaxation:
$\min_{Y\in\mathbb{R}^{n_{1}\times n_{2}}}\quad\left\{\|Y\|_{*}\ :\ Y_{ij}=M_{ij},\ \forall\,(i,j)\in\Omega\right\}.$
The above problem can be rephrased as the following SDP:
$\min_{X\in\mathbb{S}^{n_{1}+n_{2}}}\Bigl{\{}\,\frac{1}{2}\mathrm{tr}(X):X=\begin{pmatrix}W_{1}&Y\\ Y^{\top}&W_{2}\end{pmatrix}\succeq 0\,,\quad Y_{i,j}=M_{i,j}\,,\;\forall\,(i,j)\in\Omega).\Bigr{\}}$ (16)
Matrix completion instances are generated randomly, using the following procedure: Given $r\leq n_{1}\leq n_{2}$, the hidden solution matrix $M$ is the product $UV^{\top}$, where the matrices $U\in\mathbb{R}^{n_{1}\times r}$ and $V\in\mathbb{R}^{n_{2}\times r}$ have independent standard Gaussian random variables as entries. Afterward, $m$ independent and uniformly random entries from $M$ are taken, where $m=\lceil\gamma\ r(n_{1}+n_{2}-r)\rceil$ and $\gamma=r\log(n_{1}+n_{2})$ is the oversampling ratio.
Problem Instance Runtime (seconds) / rank
Size $(n_{1},n_{2};m)$ $r$ cuHALLaR cuLoRADS Ratio
3,000, 7,000; 828,931 3 0.77/3 123.71/19 160.662
3,000, 7,000; 828,931 3 0.7/3 11.42/19 16.314
3,000, 7,000; 2,302,586 5 1.28/5 3.19/19 2.492
3,000, 7,000; 2,302,586 5 1.09/5 2.26/19 2.073
10,000, 21,623; 2,948,996 3 1.22/3 5.14/21 4.213
10,000, 21,623; 8,191,654 5 1.95/5 4.71/21 2.415
25,000, 50,000; 3,367,574 2 1.2/2 10.77/23 8.975
25,000, 50,000; 7,577,040 3 1.82/3 21.31/23 11.708
30,000, 70,000; 4,605,171 2 1.63/2 17.02/24 10.441
30,000, 70,000; 10,361,633 3 2.57/3 18.45/24 7.178
50,000, 100,000; 7,151,035 2 1.89/2 23.67/24 12.523
50,000, 100,000; 16,089,828 3 3.68/3 27.95/24 7.595
80,000, 120,000; 9,764,859 2 2.45/2 20.66/25 8.432
80,000, 120,000; 21,970,931 3 5.02/3 32.28/25 6.430
80,000, 120,000; 7,996,791 3 2.80/3 17.94/25 6.407
120,000, 180,000; 11,996,886 3 3.74/3 36.47/25 9.751
160,000, 240,000; 15,996,836 3 4.90/3 52.61/25 10.736
240,000, 360,000; 23,996,846 3 7.03/3 86.80/25 12.347
Table 2: cuHALLaR vs cuLoRADS runtimes for the matrix completion problem. Seed was 0 for all instances. Tolerance is set at $10^{-5}$.
Table 2 provides a direct comparison between cuHALLaR and cuLoRADS on matrix completion problems, revealing cuHALLaR’s substantial performance advantages. Our GPU implementation consistently outperforms cuLoRADS across all problem instances with speedups ranging from 2x to over 160x. Most notably, cuHALLaR solves these problems at a significantly lower rank, demonstrating its superior ability to find optimal low-rank solutions efficiently. This comparison highlights an advantage of our approach: cuLoRADS struggles with larger problems due to its reliance on the SDPA format and less efficient memory handling, while cuHALLaR’s operator-based formulation enables it to tackle much larger instances without encountering the same memory limitations.
Problem Instance Runtime (seconds) / rank
Size $(n_{1},n_{2};m)$ $r$ HALLaR cuHALLaR Ratio
3,000, 7,000; 828,931 3 7.63/3 0.77/3 9.91
3,000, 7,000; 828,931 3 7.52/3 0.70/3 10.74
3,000, 7,000; 2,302,586 5 34.52/5 1.28/5 26.97
3,000, 7,000; 2,302,586 5 34.42/5 1.09/5 31.58
10,000, 21,623; 2,948,996 3 38.88/3 1.22/3 31.87
10,000, 21,623; 2,948,996 3 34.80/3 1.11/3 31.35
10,000, 21,623; 8,191,654 5 141.17/5 2.05/5 68.86
10,000, 21,623; 8,191,654 5 140.55/5 1.95/5 72.08
25,000, 50,000; 3,367,574 2 43.52/2 1.20/2 36.27
25,000, 50,000; 7,577,040 3 79.93/3 1.82/3 43.92
30,000, 70,000; 4,605,171 2 60.57/2 1.63/2 37.16
30,000, 70,000; 10,361,633 3 145.76/3 2.57/3 56.72
50,000, 100,000; 7,151,035 2 121.63/2 1.89/2 64.35
50,000, 100,000; 16,089,828 3 242.00/3 3.68/3 65.76
80,000, 120,000; 9,764,859 2 237.38/2 2.45/2 96.89
80,000, 120,000; 21,970,931 3 490.07/3 5.02/3 97.62
80,000, 120,000; 7,996,791 3 217.97/3 2.80/3 77.85
120,000, 180,000; 11,996,886 3 379.99/3 3.74/3 101.60
160,000, 240,000; 15,996,836 3 430.87/3 4.90/3 87.93
240,000, 360,000; 23,996,846 3 834.17/3 7.03/3 118.66
320,000, 480,000; 31,996,932 3 1,152.68/3 9.60/3 120.07
400,000, 600,000; 39,996,783 3 1,587.08/3 13.05/3 121.62
480,000, 720,000; 47,996,814 3 2,299.54/4 16.06/3 143.18
560,000, 840,000; 55,996,812 3 2,126.11/4 18.23/3 116.63
640,000, 960,000; 63,996,921 3 2,659.69/3 20.53/3 129.55
720,000, 1,080,000; 71,996,873 3 2,904.02/3 26.90/3 107.96
800,000, 1,200,000; 79,996,838 3 2,833.93/3 27.05/3 104.77
1,600,000, 2,400,000; 159,996,869 3 8,364.62/3 60.68/3 137.85
2,400,000, 3,600,000; 239,996,782 3 Timed Out 100.17/3 N/A
3,200,000, 4,800,000; 319,996,871 3 Timed Out 142.27/3 N/A
Table 3: Performance comparison between HALLaR (CPU) and cuHALLaR for matrix completion problems. A relative tolerance of $\epsilon=10^{-5}$ is used throughout. Ratio is CPU Time / GPU Time. Time limit was set at 12 hours.
Table 3 highlights the speedup achieved by cuHALLaR over its CPU counterpart. The GPU implementation consistently delivers 10-140x faster solution times across standard problem sizes while maintaining identical solution quality. Below the midrule, we showcase truly large-scale instances with dimensions from $(n_{1},n_{2})=(320,000,480,000)$ up to $(3,200,000,4,800,000)$. For these instances, cuHALLaR achieves the most remarkable speedups of 104-143x. Most significantly, for problems with dimensions $(n_{1},n_{2})=(2,400,000,3,600,000)$ and higher, the CPU implementation times out after 12 hours while cuHALLaR solves them in under 2.5 minutes. These “extra-large” instances couldn’t be evaluated on cuLoRADS due to its reliance on the SDPA format, which becomes prohibitively memory-intensive for dense problems of this scale. This demonstrates not only the efficiency of cuHALLaR, but also its ability to handle problem sizes previously deemed computationally intractable and would have taken hours on the original CPU HALLaR implementation.
Note that the experimental results in Table 3 are divided into two distinct sets of problem instances. The first set showcases standard problem dimensions where both CPU and GPU implementations can successfully operate, allowing for direct performance comparison. Below the midrule, we present a set of significantly larger-scale problems with dimensions ranging from $(n_{1},n_{2})=(320,000,480,000)$ up to $(3,200,000,4,800,000)$.
4.3 Experiments for maximum stable set
Given a graph $G=([n],E)$, the maximum stable set problem consists of finding a subset of vertices of largest cardinality such that no two vertices are connected by an edge. Lovász introduced a constant, the $\vartheta$-function, which upper bounds the value of the maximum stable set. The $\vartheta$-function is the value of the SDP
$\max\quad\{{e}{e}^{\top}\bullet X\;:\;X_{ij}=0,\;ij\in E,\;\mathrm{Tr}(X)=1,\;X\succeq 0,\quad X\in\mathbb{S}^{n}\}$ (17)
where $e=(1,1,\dots,1)\in\mathbb{R}^{n}$ is the all ones vector. It was shown in [28] that the $\vartheta$-function agrees exactly with the stable set number for perfect graphs.
Problem Instance Runtime (seconds) / rank
Problem Size $(n;m)$ Graph Dataset HALLaR cuHALLaR Ratio
10,937; 75,488 wing_nodal DIMACS10 3,826.34/130 246.71/127 15.51
16,384; 49,122 delaunay_n14 DIMACS10 643.90/36 102.43/34 6.29
16,386; 49,152 fe-sphere DIMACS10 21.74/3 17.82/3 1.22
22,499; 43,858 cs4 DIMACS10 749.95/15 588.46/115 1.27
25,016; 62,063 hi2010 DIMACS10 2,607.24/15 132.85/21 19.63
25,181; 62,875 ri2010 DIMACS10 1,911.10/17 1,951.28/190 0.98
32,580; 77,799 vt2010 DIMACS10 2,952.84/18 196.59/60 15.02
48,837; 117,275 nh2010 DIMACS10 7,694.06/20 2,186.74/172 3.52
24,300; 34,992 aug3d GHS_indef 30.10/1 2.08/1 14.47
32,430; 54,397 ia-email-EU Network Repo 598.76/5 52.86/4 11.33
11,806; 32,730 Oregon-2 SNAP 1,581.11/23 467.75/55 3.38
21,363; 91,286 ca-CondMat SNAP 8,521.04/59 348.59/77 24.44
31,379; 65,910 as-caida_G_001 SNAP 2,125.41/8 428.41/27 4.96
26,518; 65,369 p2p-Gnutella24 SNAP 312.75/4 55.99/10 5.59
22,687; 54,705 p2p-Gnutella25 SNAP 242.42/4 88.53/13 2.74
36,682; 88,328 p2p-Gnutella30 SNAP 506.26/5 61.84/10 8.19
62,586; 147,892 p2p-Gnutella31 SNAP 1,484.71/5 245.33/29 6.05
49,152; 69,632 cca AG-Monien 63.09/2 12.30/2 5.13
49,152; 73,728 ccc AG-Monien 14.52/2 20.62/2 0.70
49,152; 98,304 bfly AG-Monien 15.72/2 10.35/2 1.52
16,384; 32,765 debr_G_12 AG-Monien 250.40/10 189.57/10 1.32
32,768; 65,533 debr_G_13 AG-Monien 471.80/9 638.54/13 0.74
65,536; 131,069 debr_G_14 AG-Monien 474.22/9 29.51/9 16.07
131,072; 262,141 debr_G_15 AG-Monien 521.88/10 28.07/10 18.59
262,144; 524,285 debr_G_16 AG-Monien 1,333.31/12 84.28/13 15.82
524,288; 1,048,573 debr_G_17 AG-Monien 6,437.05/12 171.81/13 37.47
1,048,576; 2,097,149 debr_G_18 AG-Monien 16,176.30/13 119.50/13 135.37
Table 4: Runtimes (in seconds) for the Maximum stable set problem. A relative tolerance of $\epsilon=10^{-5}$ is set.
The results in Table 4 demonstrate a strong scaling advantage for GPU implementation as problem size increases. For the largest instances, cuHALLaR achieves a remarkable 135x speedup over CPU implementation. Performance patterns vary across graph families: DIMACS10 instances show substantial speedups up to 19.6x, SNAP graphs deliver consistent 3-24x improvements, and AG-Monien graphs exhibit dramatic acceleration of 16-135x for larger instances. Some smaller or sparser graphs show modest or even negative speedups, highlighting that GPU acceleration benefits depend on sufficient computational complexity to amortize parallelization overhead. These findings reconfirm that GPU acceleration becomes increasingly advantageous as problem dimensions grow, enabling cuHALLaR to efficiently solve million-scale instances that would be computationally prohibitive on CPU alone.
Problem Instance HALLaR cuHALLaR cuLoRADS HALLaR/cuHALLaR cuLoRADS/cuHALLaR
Graph($n$; $|E|$) Time/Rank Time/Rank Time/Rank Ratio Ratio
$H_{10,2}$(1,024; 5,120) 0.153/2 1.735/2 7.83/14 0.088 4.51
$H_{11,2}$(2,048; 11,264) 0.260/2 1.627/2 3.95/16 0.160 2.43
$H_{12,2}$(4,096; 24,576) 0.606/2 2.033/2 4.47/17 0.298 2.20
$H_{13,2}$(8,192; 53,248) 1.852/2 3.106/2 6.78/19 0.596 2.18
$H_{14,2}$(16,384; 114,688) 3.843/2 3.170/2 79.04/20 1.212 24.93
$H_{15,2}$(32,768; 245,760) 8.240/2 3.223/2 OOM* 2.557 N/A
$H_{16,2}$(65,536; 524,288) 17.235/2 4.457/2 OOM* 3.869 N/A
$H_{17,2}$(131,072; 1,114,112) 33.342/2 3.904/2 OOM* 8.541 N/A
$H_{18,2}$(262,144; 2,359,296) 46.436/2 2.627/2 OOM* 17.681 N/A
$H_{19,2}$(524,288; 4,980,736) 105.408/2 5.281/2 OOM* 19.958 N/A
$H_{20,2}$(1,048,576; 10,485,760) 249.701/2 8.109/2 OOM* 30.786 N/A
$H_{21,2}$(2,097,152; 22,020,096) 592.696/2 16.966/2 OOM* 34.934 N/A
$H_{22,2}$(4,194,304; 46,137,344) 1,702.255/2 37.145/2 OOM* 45.826 N/A
$H_{23,2}$(8,388,608; 96,468,992) 4,205.741/2 83.905/2 OOM* 50.125 N/A
Table 5: Runtimes (in seconds) for the Maximum Stable Set problem on Hamming graphs. A relative tolerance of $\epsilon=10^{-5}$ is set.
Table 5 shows the significant performance advantage of GPU acceleration for larger-scale Maximum Stable Set problems on Hamming graphs. For small Hamming graphs, such as $H_{10,2}$ through $H_{13,2}$, HALLaR outperforms the cuHALLaR by factors of 0.088x to 0.596x, reflecting GPU parallelization overhead for modest problem sizes and the fact that HALLaR is not designed for small problems. A clear performance transition occurs at $H_{14,2}$, where cuHALLaR begins demonstrating superior performance with a 1.21x speedup. This advantage grows exponentially as problem size increases, reaching an impressive 50.1x speedup for $H_{23,2}$ with over 8 million vertices. All instances maintain a consistent optimal solution rank of 2, confirming the effectiveness of our low-rank approach. Notably, cuLoRADS encounters memory limitations (OOM) for problems larger than $H_{14,2}$, highlighting the superior memory efficiency of our implementation. For problems it could solve, cuLoRADS shows significantly longer runtimes and higher ranks than both our GPU and CPU implementations.
The GSET benchmark [75] highlights how GPU performance shows a strong correlation with problem structure: for dense graphs with high-rank solutions such as G58, G59, and G64, GPU acceleration yields 4-5x speedups. However, for sparse graphs with very low-rank solutions such as G11, G12, and G32, CPU implementation performs 30-100x faster. This suggests that GPU acceleration is most beneficial when the computational workload involves more complex matrix operations that can effectively utilize parallel processing. Interestingly, the solution ranks sometimes differ between implementations, with GPU occasionally finding different rank solutions than CPU. The performance correlation appears more strongly tied to edge density than vertex count, with the highest speedups observed on instances having both moderate-to-large size and high edge count. These findings indicate that problem-specific characteristics, rather than simply problem size, determine when GPU acceleration provides the most significant performance advantages.
Problem Instance HALLaR cuHALLaR Ratio
G1(800; 19,176) 98.25/98 45.19/98 2.17
G10(800; 19,176) 80.51/97 40.52/97 1.99
G11(800; 1,600) 0.14/2 4.23/2 0.03
G12(800; 1,600) 0.09/2 2.59/2 0.03
G14(800; 4,694) 45.37/73 214.73/74 0.21
G20(800; 4,672) 129.73/124 359.40/166 0.36
G43(1000; 9,990) 34.20/60 35.84/60 0.95
G51(1,000; 5,909) 167.03/141 273.84/128 0.61
G23(2,000; 19,990) 94.81/76 52.59/76 1.80
G31(2,000; 19,990) 122.07/77 58.23/76 2.10
G32(2,000; 4,000) 0.35/2 25.34/3 0.01
G34(2,000; 4,000) 2.00/2 25.29/3 0.08
G35(2,000; 11,778) 397.32/133 254.68/108 1.56
G41(2,000; 11,785) 343.84/143 286.82/139 1.20
G48(3,000; 6,000) 2.29/2 18.95/2 0.12
G55(5,000; 12,498) 152.03/45 81.11/45 1.87
G56(5,000; 12,498) 151.99/45 80.83/45 1.88
G57(5,000; 10,000) 8.71/2 42.55/2 0.20
G58(5,000; 29,570) 2,191.49/112 504.55/130 4.35
G59(5,000; 29,570) 2,167.27/111 539.28/121 4.02
G60(7,000; 17,148) 262.35/50 96.89/50 2.71
G62(7,000; 14,000) 7.80/3 25.18/3 0.31
G64(7,000; 41,459) 1,719.79/51 334.12/51 5.15
G66(9,000; 18,000) 11.31/3 27.91/3 0.41
G67(10,000; 20,000) 2.39/2 5.83/2 0.41
G72(10,000; 20,000) 2.41/2 5.81/2 0.41
G77(14,000; 28,000) 27.34/3 11.81/2 2.31
G81(20,000; 40,000) 53.85/3 64.20/3 0.84
Table 6: Runtimes (in seconds) for the Maximum Stable Set problem on GSET instances. Tolerances are set to $10^{-5}$.
4.4 Experiments for phase retrieval
Given $m$ pairs $\{(a_{i},b_{i})\}_{i=1}^{m}\subseteq\mathbb{C}^{n}\times\mathbb{R}_{+}$, consider the problem of finding a vector $x\in\mathbb{C}^{n}$ such that
$|\langle a_{i},x\rangle|^{2}=b_{i},\quad i=1,\dots,m.$
In other words, the goal is to retrieve $x$ from the magnitude of $m$ linear measurements. By creating the complex Hermitian matrix $X=xx^{H}$, this problem can be approached by solving the complex-valued SDP relaxation
$\min_{X}\quad\left\{\mathrm{tr}(X)\quad:\quad\langle a_{i}a_{i}^{H},X\rangle=b_{i},\quad X\succeq 0,\quad X\in\mathbb{S}^{n}(\mathbb{C})\right\},$
[…]
Problem Instance HALLaR cuHALLaR Ratio
10,000; 120,000 53.642/5 40.866/5 1.31
10,000; 120,000 18.384/2 5.529/4 3.32
10,000; 120,000 23.644/4 4.347/3 5.44
10,000; 120,000 14.957/3 3.462/3 4.32
31,623; 379,476 149.416 29.016/5 5.15
31,623; 379,476 119.499/4 30.964/4 3.86
31,623; 379,476 153.16/5 29.279/5 5.23
31,623; 379,476 127.10/4 30.964/5 4.11
100,000; 1,200,000 303.986/5 10.401/4 29.24
100,000; 1,200,000 330.00/6 8.641/4 38.18
100,000; 1,200,000 434.10/5 9.279/4 46.78
100,000; 1,200,000 380.05/6 9.334/4 40.75
316,228; 3,704,736 600.62/2 29.016/6 20.70
316,228; 3,704,736 699.193/2 30.964/6 22.58
316,228; 3,704,736 507.60/2 33.738/7 15.05
316,228; 3,704,736 620.56/2 34.124/8 18.19
3,162,278; 37,947,336 Timed Out 583.782/14 N/A
Table 7: Runtimes (in seconds) for the Phase Retrieval problem. A relative tolerance of $\epsilon=10^{-5}$ is set and a time limit of 43200 seconds (12 hours) is given.
In Table 7 we observe significant performance improvements with cuHALLaR compared to the CPU HALLaR as problem sizes increase. For medium-sized instances with $n=31,623$, cuHALLaR achieves consistent speedups of 3.86-5.23x, completing calculations in approximately 29-31 seconds versus 119-153 seconds on CPU. For larger instances with $n=100,000$, the speedup increases dramatically to 29.24-46.78x, with cuHALLaR solving problems in just 8.6-10.4 seconds compared to 304-434 seconds on CPU. The largest reported instances of $n=316,228$ demonstrate impressive speedups of 15.05-22.58x, with GPU solution times around 29-34 seconds compared to 508-699 seconds on CPU. Most notably, cuHALLaR successfully solves the massive instance with $n=3,162,278$ in under 10 minutes, while the CPU implementation times out after 12 hours. It’s important to highlight that cuLoRADS was unable to handle these phase retrieval problems at all due to its reliance on the SDPA format, which becomes prohibitively expensive in both memory and computational requirements for problems of this scale. This further demonstrates the superior memory efficiency and scalability of our GPU implementation compared to both CPU-based methods and other GPU-accelerated SDP solvers.
5 Concluding Remarks
In this paper, we address the fundamental challenge of developing scalable and efficient solvers for large-scale semidefinite programming problems. We introduce cuHALLaR, a GPU-accelerated implementation of the hybrid low-rank augmented Lagrangian method originally proposed in [47]. Our approach exploits the parallel computing capabilities of modern GPUs to overcome computational bottlenecks inherent in handling massive SDP instances, while maintaining the advantages of low-rank factorization-based methods.
Through extensive numerical experiments across three problem domains, we demonstrate that cuHALLaR achieves remarkable performance improvements over both CPU-based implementations and competing GPU-accelerated solvers. For matrix completion problems, cuHALLaR consistently outperforms the state-of-the-art GPU solver cuLoRADS by factors of 2x to 25x, while finding solutions with significantly lower ranks. Most impressively, cuHALLaR successfully solves instances with more than 300 million constraints in under 3 minutes, while the CPU implementation, HALLaR, fails to converge within 12 hours.
For maximum stable set problems, our results reveal that cuHALLaR’s performance advantage scales dramatically with problem size and graph density. We achieve speedups of 50x for the largest Hamming graph problems and 135x for million-node AG-Monien instances. Similarly, for phase retrieval problems, cuHALLaR delivers speed improvements exceeding 74x, solving massive instances with over 3 million dimensions in under 10 minutes—problems that remain intractable for CPU-based approaches.
Future work. While cuHALLaR demonstrates exceptional performance across the problems tested, certain structural characteristics of SDPs may impact its efficiency. Problems with inherently high-rank solutions or those requiring extremely high precision may benefit less from our approach. Additionally, our current implementation focuses on single-GPU execution and double-precision arithmetic. Future research directions include developing optimized multi-GPU implementations in C/C++ to further enhance scalability, exploring mixed-precision computation techniques to balance accuracy and performance, and extending our methodology to a broader class of SDP problems. These developments would further advance the approaches introduced in [47] and establish GPU-accelerated low-rank methods as the standard for solving large-scale SDPs in practical applications.
Acknowledgments
This research was supported in part through research cyberinfrastructure resources and services provided by the Partnership for an Advanced Computing Environment (PACE) at Georgia Tech, Atlanta, Georgia, USA.
Appendix A ADAP-AIPP
This section presents the ADAP-AIPP method first developed in [47, 60]. HALLaR uses ADAP-AIPP in its step 4 to find an approximate stationary point (according to the criteria in (6)) of a nonconvex problem of the form
$\min_{U}\quad\left\{g(U)=\mathcal{L}_{\beta}(UU^{\top};p)\quad:\quad U\in B_{1}^{s}\right\}$ (18)
where $\mathcal{L}_{\beta}(X;p)$ is as in (5) and $B_{1}^{s}$ is as in (2).
Given an initial point $\underline{W}\in B_{1}^{s}$ and a tolerance $\rho>0$, the goal of ADAP-AIPP is to find a a triple $(\overline{W};\overline{R},\rho)$ that satisfies
$\overline{R}\in\nabla g(\overline{W})+N_{B^{s}_{1}}(\overline{W}),\quad\|\overline{R}\|_{F}\leq\rho,\quad g(\overline{W})\leq g(\underline{W}).$ (19)
The ADAP-AIPP method is now formally presented.
ADAP-AIPP Method
Universal Parameters: $\sigma\in(0,1/2)$ and $\chi\in(0,1)$.
Input: a function $g$ as in (18), an initial point $\underline{W}\in B_{1}^{s}$, an initial prox stepsize $\lambda_{0}>0$, and a tolerance $\rho>0$.
  • 0.
    set $W_{0}=\underline{W}$, $j=1$, and
    $\lambda=\lambda_{0},\quad\bar{M}_{0}=1;$ (20)
  • 1.
    choose $\underline{M}_{j}\in[1,\bar{M}_{j-1}]$ and call the ADAP-FISTA method in Subsection A.1 with universal input $(\sigma,\chi)$ and inputs
    $x_{0}$ $=W_{j-1},\quad(\mu,L_{0})=(1/2,\underline{M}_{j}),$ (21)
    $\psi_{s}$ $={\lambda}g+\frac{1}{2}\|\cdot-W_{j-1}\|_{F}^{2},\quad\psi_{n}={\lambda}\delta_{B^{s}_{1}};$ (22)
  • 2.
    if ADAP-FISTA fails or its output $(W,V,L)$ (if it succeeds) does not satisfy the inequality
    $\lambda g(W_{j-1})-\left[\lambda g(W)+\frac{1}{2}\|W-W_{j-1}\|_{F}^{2}\right]\geq V\bullet(W_{j-1}-W),$ (23)
    then set ${\lambda}={\lambda}/2$ and go to step $1$; else, set $({\lambda}_{j},\bar{M}_{j})=({\lambda},L)$, $(W_{j},V_{j})=(W,V)$, and
    $R_{j}:=\frac{V_{j}+W_{j-1}-W_{j}}{\lambda_{j}}$ (24)
    and go to step 3;
  • 3.
    if $\|R_{j}\|_{F}\leq\rho$, then stop with success and output $(\overline{W},\overline{R})=(W_{j},R_{j})$; else, go to step 4;
  • 4.
    set $j\leftarrow j+1$ and go to step 1.
Several remarks about ADAP-AIPP are now given.
  1. 1.
    It is shown in [47] that ADAP-AIPP is able to find a triple that satisfies (19) in $\mathcal{O}(1/\rho^{2})$ number of iterations.
  2. 2.
    ADAP-AIPP is a double looped algorithm, whose outer iterations are indexed by $j$. During its $j$-th iteration, ADAP-AIPP attempts to solve a proximal subproblem of the form $\min_{U\in B_{1}^{s}}\lambda g(U)+0.5\|\cdot-W_{j-1}\|^{2}_{F}$, where $\lambda$ is a prox stepsize and $W_{j-1}$ is the current prox center. If a proximal subproblem cannot suitably be solved, ADAP-AIPP reduces $\lambda$ and tries to solve the new subproblem.
  3. 3.
    ADAP-AIPP uses the ADAP-FISTA method, which is presented in the next Subsection A.1, to attempt to solve its proximal subproblems. For a more comprehensive treatment of ADAP-FISTA, see Appendix B in [47].
A.1 ADAP-FISTA
This subsection presents the ADAP-FISTA method that ADAP-AIPP invokes to solve proximal subproblems of the form
$\min_{U\in B_{1}^{s}}\psi_{s}(U):=\lambda g(U)+0.5\|\cdot-W\|^{2}_{F}$ (25)
where $g$ is as in (18). The ADAP-FISTA method is now formally presented.
ADAP-FISTA
Universal Parameters: $\sigma>0$ and $\chi\in(0,1)$.
Input: a function $\psi_{s}$ as in (25), an initial point $x_{0}\in B_{1}^{s}$ and scalars $\mu>0$ and $L_{0}>\mu$.
  • 0.
    set $y_{0}=x_{0}$, $A_{0}=0$, $\tau_{0}=1$, and $i=0$;
  • 1.
    set $L_{i+1}=L_{i}$;
  • 2.
    compute
    $a_{i}$ $=\frac{\tau_{i}+\sqrt{\tau_{i}^{2}+4\,\tau_{i}A_{i}\,(L_{i+1}-\mu)}}{2\,(L_{i+1}-\mu)},\quad\tilde{x}_{i}=\frac{A_{i}\,y_{i}+a_{i}\,x_{i}}{A_{i}+a_{i}},$ (26)
    $y_{i+1}$ $\;\;=\;\;\arg\min_{u\in B_{1}^{s}}\Bigl{\{}\ell_{\psi_{s}}(u;\tilde{x}_{i})+\tfrac{L_{i+1}}{2}\|u-\tilde{x}_{i}\|^{2}\Bigr{\}};$ (27)
    if
    $\ell_{\psi_{s}}(y_{i+1};\tilde{x}_{i})+\tfrac{(1-\chi)\,L_{i+1}}{4}\,\|y_{i+1}-\tilde{x}_{i}\|^{2}\;\geq\;\psi_{s}(y_{i+1}),$ (28)
    go to step 3; else set $L_{i+1}\leftarrow 2\,L_{i+1}$ and repeat step 2;
  • 3.
    update
    $A_{i+1}$ $=A_{i}+a_{i},\qquad\tau_{i+1}=\tau_{i}+a_{i}\,\mu,$ (29)
    $s_{i+1}$ $=(L_{i+1}-\mu)\,(\tilde{x}_{i}-y_{i+1}),$ (30)
    $x_{i+1}$ $=\frac{1}{\tau_{i+1}}\bigl{[}\mu\,a_{i}\,y_{i+1}+\tau_{i}\,x_{i}-a_{i}\,s_{i+1}\bigr{]};$ (31)
  • 4.
    if
    $\|y_{i+1}-x_{0}\|^{2}\;\geq\;\chi\,A_{i+1}\,L_{i+1}\,\|y_{i+1}-\tilde{x}_{i}\|^{2},$ (32)
    then go to step 5; otherwise stop with failure;
  • 5.
    Compute
    $v_{i+1}=\nabla\psi_{s}(y_{i+1})-\nabla\psi_{s}(\tilde{x}_{i})+L_{i+1}(\tilde{x}_{i}-y_{i+1}).$ (33)
    If the inequality
    $\|v_{i+1}\|\leq\sigma\|y_{i+1}-x_{0}\|$ (34)
    holds then stop with success and output $(y,v,L):=(y_{i+1},v_{i+1},L_{i+1})$; otherwise, $i\leftarrow i+1$ and go to step 1.
Several remarks about ADAP-FISTA are now given. ADAP-FISTA is either able to successfully find a suitable solution of (25) or is unable to do so in at most
$\mathcal{O}\left(\sqrt{L_{\psi_{s}}}\right)$
number of iterations, where $L_{\psi_{s}}$ is the Lipschitz constant of the gradient of the objective in (25). If the objective in (25) is strongly convex, then ADAP-FISTA always succeeds in solving (25). For more technical details on the type of solution that ADAP-FISTA aims to find, see Appendix B in [47].